OpenStudy (anonymous):

Which statement is true? Answer a) Any two isosceles triangles are similar. b) Any two triangles are similar. c) Any two right triangles are similar. d) Any two equilateral triangles are similar.

OpenStudy (anonymous):

Try to think of nonsimilar triangles that meet the criteria in each one.

OpenStudy (anonymous):

I got equilateral, but I wasn't sure.

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

That's right. Here's why. You have two triangles. One has sides a, b, and c. The second has sides d, e, and f. If they are similar, then a/d = b/e = c/f. Now if you have equilateral triangles, then b=a and c=a. Also, e=d and f=d. So a/d = a/d = a/d.

OpenStudy (amistre64):

i couldnt recall the diff between iso and equi lol

OpenStudy (anonymous):

ahahaha I relate. When I did my student teaching in a high school geometry class, I had to relearn like 70% of the vocabulary.

hero (hero):

That's bizarre, especially for you amistre